# Geometric Transformations on Null Curves in the Anti-de Sitter $3$-Space

### v1

@@ 3.2. Case $\chi=0$
Let $\gamma:J\subseteq\mathbb{R}\longrightarrow{\rm AdS}$ be a null curve with bending $\kappa$ and $(F_{+},F_{-})$ be the spinor frame field along it. A null curve $\widetilde{\gamma}:J\subseteq\mathbb{R}\longrightarrow{\rm AdS}$ is a $\mathcal{T}$-transform of $\gamma$ for $\chi=0$ if and only if
(3.13) $\widetilde{\gamma}=\pm F_{+}\begin{pmatrix}\tanh(\xi)&-{\rm csch}(\xi){\rm sech}(\xi)f\\ 0&\coth(\xi)\end{pmatrix}F_{-}^{-1}\,,$
where $\xi\neq 0$ is a constant and $f:J\subseteq\mathbb{R}\longrightarrow\mathbb{R}$ is a solution of the Riccati equation
(3.14) $f^{\prime}+f^{2}=\kappa+\cosh(2\xi)\,.$
@@
(3.15) $\widetilde{\kappa}=-\kappa+2f^{2}-2\cosh(2\xi)\,.$
@@
Let $\gamma$ be a null curve with bending $\kappa$ and assume that $\widetilde{\gamma}$ is a $\mathcal{T}$-transform of $\gamma$ for $\chi=0$. The proof of Proposition 3.5 shows that $f_{+}=f_{-}$ and $02+c_{+}^{2}-c_{-}^{2}=$ 0. From the latter there must exist a nonzero constant $\xi$ such that $c_{+}=\sqrt{2}\sinh(\xi)$ and $c_{-}=\pm\sqrt{2}\cosh(\xi)$. Moreover, the function $f_{+}=f_{-}$ is a solution of the Riccati equation (3.5). For simplicity, we simply put $f=f_{\pm}$. The Riccati equation (3.5) now reads
@@
$=$ ±$\pm F_{+}\begin{pmatrix}\tanh(\xi)&-{\rm csch}(\xi){\rm sech}(\xi)f\\ 0&\coth(\xi)\end{pmatrix}F_{-}^{-1}\,,$
@@
This concludes the forward implication. Conversely, suppose that $f:J\subseteq\mathbb{R}\longrightarrow\mathbb{R}$ is a solution of the Riccati equation (3.14) and consider the ${\rm SL}(2,\mathbb{R})\times{\rm SL}(2,\mathbb{R})$-valued map defined by $(\widetilde{F}_{+},\widetilde{F}_{-})$ where
(3.20) $\widetilde{F}_{+}$ $=$ $\frac{1}{\sqrt{2}\sinh(\xi)}\,F_{+}\begin{pmatrix}-f&f^{2}-2\sinh^{2}(\xi)\\ 1&-f\end{pmatrix},$
(3.23) $\widetilde{F}_{-}$ $=$ ±$\frac{\pm 1}{\sqrt{2}\cosh(\xi)}\,F_{-}\begin{pmatrix}-f&f^{2}-2\cosh^{2}(\xi)\\ 1&-f\end{pmatrix}.$
@@
Then, the map $(\widetilde{F}_{+},\widetilde{F}_{-})$ is a lift of a null curve $\widetilde{\gamma}$ to ${\rm SL}(2,\mathbb{R})\times{\rm SL}(2,\mathbb{R})$. We will first show that $(\widetilde{F}_{+},\widetilde{F}_{-})$ is indeed the spinor frame field along $\widetilde{\gamma}$.
From the spinorial Frenet-type equations (2.3) and (2.6) of $\gamma$ and the definition of $\widetilde{F}_{\pm}$ given in (3.20)-(3.23), it follows that
$\widetilde{F}_{+}^{-1}d\widetilde{F}_{+}=\begin{pmatrix}0&\widetilde{\kappa}+1\\ 1&0\end{pmatrix},\quad\quad\quad\widetilde{F}_{-}^{-1}d\widetilde{F}_{-}=\begin{pmatrix}0&\widetilde{\kappa}-1\\ 1&0\end{pmatrix},$
where $\widetilde{\kappa}$ is as in (3.15). Consequently, $(\widetilde{F}_{+},\widetilde{F}_{-})$ is the spinor frame field along $\widetilde{\gamma}$ and the function $\widetilde{\kappa}$ is the bending of $\widetilde{\gamma}$.
We now prove that $\widetilde{\gamma}$ is a $\mathcal{T}$-transform of $\gamma$. Since $(\widetilde{F}_{+},\widetilde{F}_{-})$ is the spinor frame field along $\widetilde{\gamma}$, the first column vectors of $\widetilde{F}_{+}$ and $\widetilde{F}_{-}$ are the pair $(\widetilde{\eta}_{+},\widetilde{\eta}_{-})$ of star-shaped cousins associated with $\widetilde{\gamma}$. From the expressions of $\widetilde{F}_{+}$ and $\widetilde{F}_{-}$ given in (3.20) and (3.23), respectively, we have
$\widetilde{\eta}_{+}$ $=$ $\frac{1}{\sqrt{2}\sinh(\xi)}\left(-f\eta_{+}+\eta_{+}^{\prime}\right),$
$\widetilde{\eta}_{-}$ $=$ ±$\frac{\pm 1}{\sqrt{2}\cosh(\xi)}\left(-f\eta_{-}+\eta_{-}^{\prime}\right).$
A simple computation involving ${\rm det}(\eta_{+},\eta_{+}^{\prime})=1$ and ${\rm det}(\eta_{-},\eta_{-}^{\prime})=1$, then shows that
${\rm det}(\eta_{+},\widetilde{\eta}_{+})=\frac{1}{\sqrt{2}\sinh(\xi)}\,,\quad\quad\quad{\rm det}(\eta_{-},\widetilde{\eta}_{-})=\frac{\pm 1}{\sqrt{2}\cosh(\xi)}\,.$
@@
The null curve $\widetilde{\gamma}$ given by (3.13) is called the $\mathcal{T}$-transform (for $\chi=0$) of $\gamma$ with parameter $\xi$ and transforming function $f$. We denote it by $\mathcal{T}_{\xi,f}(\gamma)$.
@@
Refer to caption   Refer to caption
Figure 4. Left: The $\mathcal{T}$-transform (for $\chi=0$) $\widetilde{\gamma}=\mathcal{T}_{\xi,f}(\gamma)$ of the null curve $\gamma$ with constant bending $\kappa_{7,3}$ (see Figure 1). The parameter $\xi\neq 0$ of the $\mathcal{T}$-transform is $\xi=1.01$ and the transforming function $f$ is the solution of (3.14) with initial condition $0.1$. Right: The associated pair of star-shaped cousins $(\widetilde{\eta}_{+},\widetilde{\eta}_{-})$ in plain blue and red, respectively. The curves dotted in yellow are the corresponding $\mathcal{T}$-transforms, namely, $\eta_{+}$ and $\eta_{-}$ (cf. Figures 2 and 3).
In the next result we prove, using standard arguments, a permutability theorem for the $\mathcal{T}$-transform on null curves in ${\rm AdS}$.
@@
Let $\gamma:J\subseteq\mathbb{R}\longrightarrow{\rm AdS}$ be a null curve with bending $\kappa$ and consider two $\mathcal{T}$-transforms of $\gamma$, namely, $\gamma_{1}=\mathcal{T}_{\xi_{1},f_{1}}(\gamma)$ and $\gamma_{2}=\mathcal{T}_{\xi_{2},f_{2}}(\gamma)$ with parameters $\xi_{1}\neq\xi_{2}$ and transforming functions $f_{1}\neq f_{2}$, respectively. Then, the functions
$f_{2\to 1}$ $=$ $-f_{1}+\frac{\cosh(2\xi_{1})-\cosh(2\xi_{2})}{f_{1}-f_{2}}\,,$
$f_{1\to 2}$ $=$ $-f_{2}+\frac{\cosh(2\xi_{1})-\cosh(2\xi_{2})}{f_{1}-f_{2}}\,,$
@@
$f_{2\to 1}^{\prime}+f_{2\to 1}^{2}=\kappa_{1}+\cosh(2\xi_{2})\,,\quad\quad\quad f_{1\to 2}^{\prime}+f_{1\to 2}^{2}=\kappa_{2}+\cosh(2\xi_{1})\,,$
@@
(3.24) $\mathcal{T}_{\xi_{2},f_{2\to 1}}\left(\mathcal{T}_{\xi_{1},f_{1}}(\gamma)\right)=\mathcal{T}_{\xi_{1},f_{1\to 2}}\left(\mathcal{T}_{\xi_{2},f_{2}}(\gamma)\right),$
and the bending of $\mathcal{T}_{\xi_{2},f_{2\to 1}}\left(\mathcal{T}_{\xi_{1},f_{1}}(\gamma)\right)$ is the function defined by
$\kappa_{2\to 1}=\kappa-2\left(\cosh(2\xi_{1})-\cosh(2\xi_{2})\right)\frac{f_{1}+f_{2}}{f_{1}-f_{2}}+2\left(\frac{\cosh(2\xi_{1})-\cosh(2\xi_{2})}{f_{1}-f_{2}}\right)^{2}.$
(Of course, the bending $\kappa_{1\to 2}$ of $\mathcal{T}_{\xi_{1},f_{1\to 2}}\left(\mathcal{T}_{\xi_{2},f_{2}}(\gamma)\right)$ is, precisely, $\kappa_{1\to 2}=\kappa_{2\to 1}$.)
@@
Let $\gamma$ be a null curve with bending $\kappa$ and assume that $\gamma_{1}=\mathcal{T}_{\xi_{1},f_{1}}(\gamma)$ and $\gamma_{2}=\mathcal{T}_{\xi_{2},f_{2}}(\gamma)$ are two $\mathcal{T}$-transforms of $\gamma$. From Theorem 3.8 it follows that
$\kappa_{1}=-\kappa+2f_{1}^{2}-2\cosh(2\xi_{1})\,,\quad\quad\quad\kappa_{2}=-\kappa+2f_{2}^{2}-2\cosh(2\xi_{2})\,,$
@@
$f_{1}^{\prime}+f_{1}^{2}=\kappa+\cosh(2\xi_{1})\,,\quad\quad\quad f_{2}^{\prime}+f_{2}^{2}=\kappa+\cosh(2\xi_{2})\,,$
@@
$G_{-}(\xi,\phi)$ $=$ ±$\frac{\pm 1}{\sqrt{2}\cosh(\xi)}\begin{pmatrix}-\phi&\phi^{2}-2\cosh^{2}(\xi)\\ 1&-\phi\end{pmatrix}.$
@@
$\left(\left[F_{+}G_{+}(\xi_{1},f_{1})\right]G_{+}(\xi_{2},f_{2\to 1}),\left[F_{-}G_{-}(\xi_{1},f_{1})\right]G_{-}(\xi_{2},f_{2\to 1})\right),$
@@
$\left(\left[F_{+}G_{+}(\xi_{2},f_{2})\right]G_{+}(\xi_{1},f_{1\to 2}),\left[F_{-}G_{-}(\xi_{2},f_{2})\right]G_{-}(\xi_{1},f_{1\to 2})\right),$
@@
(3.27) $\kappa_{2\to 1}=-\kappa_{1}+2f_{2\to 1}^{2}-2\cosh(2\xi_{2})\,.$
At the same time $\gamma_{1}=\mathcal{T}_{\xi_{1},f_{1}}(\gamma)$ is a $\mathcal{T}$-transform of $\gamma$ with parameter $\xi_{1}$ and transforming function $f_{1}$. Hence, $\kappa_{1}=-\kappa+2f_{1}^{2}-2\cosh(2\xi_{1})$ which we substitute in (3.27), obtaining
@@
4. Geometric Realization of the Bäcklund Transformation for the KdV
@@
4.1. Bäcklund Transformation for the KdV
The Korteweg-De Vries (KdV) equation is the PDE given by
(4.1) $\partial_{t}\kappa+\partial_{s}^{3}\kappa-6\kappa\,\partial_{s}\kappa=0\,.$
Let $J,I\subseteq\mathbb{R}$ be two open intervals (for convenience, we assume $0\in J,I$) and consider a solution $\kappa:(s,t)\in J\times I\subseteq\mathbb{R}^{2}\longmapsto\kappa(s,t)\in\mathbb{R}$ of the KdV equation (4.1). The Bäcklund transform of $\kappa$ with spectral parameter $\lambda\in\mathbb{R}\setminus\{0\}$ and transforming function $f:J\times I\subseteq\mathbb{R}^{2}\longrightarrow\mathbb{R}$ is the function defined by
(4.2) $\widetilde{\kappa}=-\kappa+2f^{2}-2\lambda\,.$
In [14], Wahlquist and Estabrook showed that if the transforming function $f(s,t)$ is a solution of the so-called Wahlquist-Estabrook equation
(4.3) $df=\left(\kappa-f^{2}+\lambda\right)ds+\left(4\lambda[f^{2}-\lambda]-2[f^{2}+\lambda]+2\kappa^{2}+2f\partial_{s}\kappa-\partial_{s}^{2}\kappa\right)dt\,,$

### v2

@@ 3.2 Case $\boldsymbol{\chi=0}$
Let $\gamma\colon J\subseteq\mathbb{R}\longrightarrow{\rm AdS}$ be a null curve with bending $\kappa$ and $(F_{+},F_{-})$ be the spinor frame field along it. A null curve $\widetilde{\gamma}\colon J\subseteq\mathbb{R}\longrightarrow{\rm AdS}$ is a $\mathcal{T}$-transform of $\gamma$ for $\chi=0$ if and only if
$\widetilde{\gamma}=\pm F_{+}\begin{pmatrix}\tanh(\xi)&-{\rm csch}(\xi){\rm sech}(\xi)f\\ 0&\coth(\xi)\end{pmatrix}F_{-}^{-1},$ (3.10)
where $\xi\neq 0$ is a constant and $f\colon J\subseteq\mathbb{R}\longrightarrow\mathbb{R}$ is a solution of the Riccati equation
$f^{\prime}+f^{2}=\kappa+\cosh(2\xi).$ (3.11)
@@
$\widetilde{\kappa}=-\kappa+2f^{2}-2\cosh(2\xi).$ (3.12)
@@
Let $\gamma$ be a null curve with bending $\kappa$ and assume that $\widetilde{\gamma}$ is a $\mathcal{T}$-transform of $\gamma$ for $\chi=0$.
The proof of Proposition 3.5 shows that $f_{+}=f_{-}$ and $02+c_{+}^{2}-c_{-}^{2}=$ 0. From the latter, there must exist a nonzero constant $\xi$ such that $c_{+}=\sqrt{2}\sinh(\xi)$ and $c_{-}=\pm\sqrt{2}\cosh(\xi)$. Moreover, the function $f_{+}=f_{-}$ is a solution of the Riccati equation (3.5). For simplicity, we simply put $f=f_{\pm}$. The Riccati equation (3.5) now reads
@@
=±${}=\pm F_{+}\begin{pmatrix}\tanh(\xi)&-{\rm csch}(\xi){\rm sech}(\xi)f\\ 0&\coth(\xi)\end{pmatrix}F_{-}^{-1},$
@@
This concludes the forward implication.
Conversely, suppose that $f\colon J\subseteq\mathbb{R}\longrightarrow\mathbb{R}$ is a solution of the Riccati equation (3.11) and consider the ${\rm SL}(2,\mathbb{R})\times{\rm SL}(2,\mathbb{R})$-valued map defined by $\bigl{(}\widetilde{F}_{+},\widetilde{F}_{-}\bigr{)}$, where
$\widetilde{F}_{+}=\frac{1}{\sqrt{2}\sinh(\xi)}\,F_{+}\begin{pmatrix}-f&f^{2}-2\sinh^{2}(\xi)\\ 1&-f\end{pmatrix},$ (3.13)
$\widetilde{F}_{-}=\frac{\pm 1}{\sqrt{2}\cosh(\xi)}\,F_{-}\begin{pmatrix}-f&f^{2}-2\cosh^{2}(\xi)\\ 1&-f\end{pmatrix}.$ (3.14)
@@
Then, the map $\bigl{(}\widetilde{F}_{+},\widetilde{F}_{-}\bigr{)}$ is a lift of a null curve $\widetilde{\gamma}$ to ${\rm SL}(2,\mathbb{R})\times{\rm SL}(2,\mathbb{R})$. We will first show that $\bigl{(}\widetilde{F}_{+},\widetilde{F}_{-}\bigr{)}$ is indeed the spinor frame field along $\widetilde{\gamma}$.
From the spinorial Frenet-type equations (2.1) and (2.2) of $\gamma$ and the definition of $\widetilde{F}_{\pm}$ given in (3.13) and (3.14), it follows that
$\widetilde{F}_{+}^{-1}{\rm d}\widetilde{F}_{+}=\begin{pmatrix}0&\widetilde{\kappa}+1\\ 1&0\end{pmatrix},\qquad\widetilde{F}_{-}^{-1}{\rm d}\widetilde{F}_{-}=\begin{pmatrix}0&\widetilde{\kappa}-1\\ 1&0\end{pmatrix},$
where $\widetilde{\kappa}$ is as in (3.12). Consequently, $\bigl{(}\widetilde{F}_{+},\widetilde{F}_{-}\bigr{)}$ is the spinor frame field along $\widetilde{\gamma}$ and the function $\widetilde{\kappa}$ is the bending of $\widetilde{\gamma}$.
We now prove that $\widetilde{\gamma}$ is a $\mathcal{T}$-transform of $\gamma$. Since $\bigl{(}\widetilde{F}_{+},\widetilde{F}_{-}\bigr{)}$ is the spinor frame field along $\widetilde{\gamma}$, the first column vectors of $\widetilde{F}_{+}$ and $\widetilde{F}_{-}$ are the pair $(\widetilde{\eta}_{+},\widetilde{\eta}_{-})$ of star-shaped cousins associated with $\widetilde{\gamma}$. From the expressions of $\widetilde{F}_{+}$ and $\widetilde{F}_{-}$ given in (3.13) and (3.14), respectively, we have
$\widetilde{\eta}_{+}=\frac{1}{\sqrt{2}\sinh(\xi)}\bigl{(}-f\eta_{+}+\eta_{+}^{\prime}\bigr{)},\qquad\widetilde{\eta}_{-}=\frac{\pm 1}{\sqrt{2}\cosh(\xi)}\bigl{(}-f\eta_{-}+\eta_{-}^{\prime}\bigr{)}.$
A simple computation involving $\det\bigl{(}\eta_{+},\eta_{+}^{\prime}\bigr{)}=1$ and $\det\bigl{(}\eta_{-},\eta_{-}^{\prime}\bigr{)}=1$, then shows that
$\det(\eta_{+},\widetilde{\eta}_{+})=\frac{1}{\sqrt{2}\sinh(\xi)},\qquad\det(\eta_{-},\widetilde{\eta}_{-})=\frac{\pm 1}{\sqrt{2}\cosh(\xi)}.$
@@
The null curve $\widetilde{\gamma}$ given by (3.10) is called the $\mathcal{T}$-transform ((((for $\chi=0)$ of $\gamma$ with parameter $\xi$ and transforming function $f$. We denote it by $\mathcal{T}_{\xi,f}(\gamma)$.
@@
In the next result, we prove, using standard arguments, a permutability theorem for the $\mathcal{T}$-transform on null curves in ${\rm AdS}$.
@@
Let $\gamma\colon J\subseteq\mathbb{R}\longrightarrow{\rm AdS}$ be a null curve with bending $\kappa$ and consider two $\mathcal{T}$-transforms of $\gamma$, namely, $\gamma_{1}=\mathcal{T}_{\xi_{1},f_{1}}(\gamma)$ and $\gamma_{2}=\mathcal{T}_{\xi_{2},f_{2}}(\gamma)$ with parameters $\xi_{1}\neq\xi_{2}$ and transforming functions $f_{1}\neq f_{2}$, respectively. Then, the functions
$f_{2\to 1}=-f_{1}+\frac{\cosh(2\xi_{1})-\cosh(2\xi_{2})}{f_{1}-f_{2}},\qquad f_{1\to 2}=-f_{2}+\frac{\cosh(2\xi_{1})-\cosh(2\xi_{2})}{f_{1}-f_{2}},$
@@
$f_{2\to 1}^{\prime}+f_{2\to 1}^{2}=\kappa_{1}+\cosh(2\xi_{2}),\qquad f_{1\to 2}^{\prime}+f_{1\to 2}^{2}=\kappa_{2}+\cosh(2\xi_{1}),$
@@
$\mathcal{T}_{\xi_{2},f_{2\to 1}}(\mathcal{T}_{\xi_{1},f_{1}}(\gamma))=\mathcal{T}_{\xi_{1},f_{1\to 2}}(\mathcal{T}_{\xi_{2},f_{2}}(\gamma)),$
and the bending of $\mathcal{T}_{\xi_{2},f_{2\to 1}}(\mathcal{T}_{\xi_{1},f_{1}}(\gamma))$ is the function defined by
$\kappa_{2\to 1}=\kappa-2(\cosh(2\xi_{1})-\cosh(2\xi_{2}))\frac{f_{1}+f_{2}}{f_{1}-f_{2}}+2\biggl{(}\frac{\cosh(2\xi_{1})-\cosh(2\xi_{2})}{f_{1}-f_{2}}\biggr{)}^{2}.$
((((Of course, the bending $\kappa_{1\to 2}$ of $\mathcal{T}_{\xi_{1},f_{1\to 2}}(\mathcal{T}_{\xi_{2},f_{2}}(\gamma))$ is, precisely, $\kappa_{1\to 2}=\kappa_{2\to 1}$.))))
@@
Let $\gamma$ be a null curve with bending $\kappa$ and assume that $\gamma_{1}=\mathcal{T}_{\xi_{1},f_{1}}(\gamma)$ and $\gamma_{2}=\mathcal{T}_{\xi_{2},f_{2}}(\gamma)$ are two $\mathcal{T}$-transforms of $\gamma$. From Theorem 3.8, it follows that
$\kappa_{1}=-\kappa+2f_{1}^{2}-2\cosh(2\xi_{1}),\qquad\kappa_{2}=-\kappa+2f_{2}^{2}-2\cosh(2\xi_{2})$
@@
$f_{1}^{\prime}+f_{1}^{2}=\kappa+\cosh(2\xi_{1}),\qquad f_{2}^{\prime}+f_{2}^{2}=\kappa+\cosh(2\xi_{2}),$
@@
$G_{-}(\xi,\phi)=\frac{\pm 1}{\sqrt{2}\cosh(\xi)}\begin{pmatrix}-\phi&\phi^{2}-2\cosh^{2}(\xi)\\ 1&-\phi\end{pmatrix}.$
@@
$([F_{+}G_{+}(\xi_{1},f_{1})]G_{+}(\xi_{2},f_{2\to 1}),[F_{-}G_{-}(\xi_{1},f_{1})]G_{-}(\xi_{2},f_{2\to 1}))$
@@
$([F_{+}G_{+}(\xi_{2},f_{2})]G_{+}(\xi_{1},f_{1\to 2}),[F_{-}G_{-}(\xi_{2},f_{2})]G_{-}(\xi_{1},f_{1\to 2}))$
@@
$\kappa_{2\to 1}=-\kappa_{1}+2f_{2\to 1}^{2}-2\cosh(2\xi_{2}).$ (3.15)
At the same time, $\gamma_{1}=\mathcal{T}_{\xi_{1},f_{1}}(\gamma)$ is a $\mathcal{T}$-transform of $\gamma$ with parameter $\xi_{1}$ and transforming function $f_{1}$. Hence, $\kappa_{1}=-\kappa+2f_{1}^{2}-2\cosh(2\xi_{1})$ which we substitute in (3.15), obtaining
@@
4 Geometric realization of the Bäcklund transformation
for the KdV
@@ 4 Geometric realization of the Bäcklund transformation
4.1 Bäcklund transformation for the KdV
The Korteweg–de Vries (KdV) equation is the PDE given by
$\partial_{t}\kappa+\partial_{s}^{3}\kappa-6\kappa\partial_{s}\kappa=0.$ (4.1)
Let $J,I\subseteq\mathbb{R}$ be two open intervals (for convenience, we assume $0\in J,I$) and consider a solution $\kappa\colon(s,t)\in J\times I\subseteq\mathbb{R}^{2}\longmapsto\kappa(s,t)\in\mathbb{R}$ of the KdV equation (4.1). The Bäcklund transform of $\kappa$ with spectral parameter $\lambda\in\mathbb{R}\setminus\{0\}$ and transforming function $f\colon J\times I\subseteq\mathbb{R}^{2}\longrightarrow\mathbb{R}$ is the function defined by
$\widetilde{\kappa}=-\kappa+2f^{2}-2\lambda.$ (4.2)
In [14], Wahlquist and Estabrook showed that if the transforming function $f(s,t)$ is a solution of the so-called Wahlquist–Estabrook equation
${\rm d}f=\bigl{(}\kappa-f^{2}+\lambda\bigr{)}{\rm d}s+\bigl{(}4\lambda\bigl{[}f^{2}-\lambda\bigr{]}-2\bigl{[}f^{2}+\lambda\bigr{]}+2\kappa^{2}+2f\partial_{s}\kappa-\partial_{s}^{2}\kappa\bigr{)}{\rm d}t,$ (4.3)